OpenStudy (anonymous):

Identify the inside function, u = g(x),and the outside function, y = f(u). y = (x^2 − 2x + 4)^2 Can someone help me figure out (u, y) = (g(x), f(u)) =? I have no instruction on this type of problem at all, so I need walked through it ALL. Please help?!!!!?

OpenStudy (jdoe0001):

hmm well.... looks to me like a case of y = f( g(x) )

OpenStudy (jdoe0001):

\(\bf y=f({\color{brown}{ u}})\qquad {\color{brown}{ u}}=g(x)\quad thus\quad y=f({\color{brown}{ u}}) \iff y = f(\quad g(x)\quad )\)

OpenStudy (jdoe0001):

so... what do you think?

OpenStudy (jdoe0001):

well the usual suspect will be that f(x) is the one with the exponent and the g(x) is the base so \(\large \bf y ={\color{blue}{ f(\quad {\color{brown}{ g(x)}}\quad )}}= {\color{blue}{ ({\color{brown}{ x^2 - 2x + 4}})^2}}\) any ideas?

OpenStudy (anonymous):

soo it would be: x^2-2x+4, u^2?

OpenStudy (jdoe0001):

yeap g(x) and f(x) respectively

OpenStudy (jdoe0001):

\(\bf f(u)=u^2\qquad {\color{brown}{ g(x)}}=x^2 - 2x + 4 \\ \quad \\ f(\quad g(x)\quad )={\color{brown}{ g(x)}}^2\to ({\color{brown}{ x^2 - 2x + 4)}})^2\leftarrow y\)

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw